Most calculators have 2 different logs on them: log, which is based 10, and ln, which is based e. In computer science, digital computers are based on the binary numbering system which means that there are only 2 numbers available to the computer, 0 and 1. When a computer scientist needs a logarithm, he/she uses a log to the base 2. To find the log of a number to any base, we can use a conversion formula as shown here:

Using this formula, find . Round your answer to the hundredth's place.
Answer:
Show work in this space

Answer by vanetiks(6)

Answer: 13.287
Log2=log10000/Log2=4/.30102= 13.287

or
10000=10^4
log 10000=log10^4
Log (base a)(a^y)=y
log (base 10)(10^4)=4
log(10000)=4
You can write your work out either way